OpenStudy (ria23):

Evaluate the discriminant for each equation, then use it to determine the nature of the roots for each quadratic equation. a.) x^2 + 4x + 5 = 0 b.) x^2 - 4x - 5 = 0 c.) 4x^2 + 20x + 25 = 0 I know the discriminant is -b/2a but I don't understand what it means by "determine the nature" or if it wants me to solve them or not...

OpenStudy (anonymous):

discriminant is actually b^2 - 4ac if discriminant D = b^2 - 4ac > 0 , function has two roots if discriminant b^2 - 4ac = 0 , function has a double root (only one solution) if disctiminant b^2 - 4ac < 0 , function has no solutions (no roots)

OpenStudy (ria23):

Oh my gosh. I totally missed that question. Thank yhu so much for yhur help. I appreciate it. c:

OpenStudy (anonymous):

of course!